how do you solve 11x^2=27? round to the nearest hundreth
Answer by Alan3354(69443)

how do you solve 11x^2=27? round to the nearest hundreth
----------------
11x^2=27
x^2 = 27/11
x = +sqrt(27/11) = 3sqrt(3/11)
x = -sqrt(27/11) = -3sqrt(3/11)
x =~ +/- 1.57